OpenStudy (juan1857):

Find the rectangular coordinates of the point with the polar coordinates: (-2, 5pi/6)

OpenStudy (john_es):

If I am correct, polar coordinates have a radius and an angle so \[x=r\cos\theta\\ y=r\sin\theta\] You have polar coordinates \[(r,\theta)=(-2,5\pi/6)\] Then you can find the x and y.

OpenStudy (juan1857):

I was thrown off trying to find x

OpenStudy (john_es):

x=-2*cos(5pi/6) y=-2sin(5pi/6)

OpenStudy (juan1857):

so for x you would get 1.9?

OpenStudy (john_es):

I would obtain 1.732 or exactly, \[x=\sqrt{3}\]

OpenStudy (john_es):

And for y \[y=-1\]

OpenStudy (juan1857):

can you explain how you put it in the calculator

OpenStudy (john_es):

That way, \[-2\cdot \cos(5\cdot 3.1415/6)\] Remember that your calculator must be in radians.

OpenStudy (juan1857):

oh ok Thanks!